In a bolt manufacturing factory, machines A,B, and C produce 25%, 30% and 45% of the total output, respectively. Of their outputs, 7%, 6% and 4% are defective bolts, respectively. If a bolt drawn at random from the production is found to be defective, what is the probability that it was manufactured by machine C?
1
Expert's answer
2020-06-01T19:41:53-0400

Let, A1 be the event that the drawn bolt is manufactured by machine A

A2 be the event that the drawn bolt is manufactured by machine B

A3 be the event that the drawn bolt is manufactured by machine C

B be the event that the drawn bolt is defective


Now, we are given,

P(A1) = 25% = 0.25, P(A2) = 30% = 0.3, P(A3) = 45% = 0.45, P(B | A1) = 7% = 0.07, P(B | A2) = 6% = 0.06, P(B | A3) = 4% = 0.04


The required probability

= The probability that the defective bolt was manufactured by machine C

= The probability that the bolt was manufactured by machine C given it was defective

= P(A3 | B)


Here the events A1, A2 and A3 are mutually exclusive and exhaustive.


By using Bayes' theorem,


P(A3 | B) = "\frac{P(A_3).P(B |A_3)}{P(A_1).P(B |A_1)+P(A_2).P(B |A_2)+P(A_3).P(B |A_3)}"


"\frac{0.04 \times 0.45}{0.07 \times 0.25+0.06 \times 0.3+0.04 \times 0.45}"


= 0.336 (rounded to 3 decimal places)


Answer: If a bolt drawn at random from the production is found to be defective, the probability that it was manufactured by machine C is 0.336.